Find the distance of the point of intersection of the line

(x-2)/1 = (y+3)/-1 = z/3

and the plane 2x- 3y +4z+ 4=0 from the origin.


1
Expert's answer
2021-06-07T19:02:04-0400
"\frac{x-2}{1}=\frac{y+3}{-1}=\frac{z}{3}=t\\
\begin{cases}
x=t+2\\
y=-t-3\\
z=3t
\end{cases}"


Substitute in the equation of the plane


"2x-3y+4z+4=0"



"2t+4+3t+9+12t+4=0"

"17t+17=0
\implies t=-1\\"

"\begin{cases}
x=1\\
y=-2\\
z=-3
\end{cases}"

The distance of the point of intersection from the origin:


"r=\sqrt{(1)^2+(-2)^2+(-3)^2} = \sqrt{14}"